# Towards Sharp Minimax Risk Bounds for Operator Learning

### v1

@@ Abstract
We develop a minimax theory for operator learning, where the goal is to estimate an unknown operator between separable Hilbert spaces from finitely many noisy input-output samples. For uniformly bounded Lipschitz operators, we prove information-theoretic lower bounds together with matching or near-matching upper bounds, covering both fixed and random designs under Hilbert-valued Gaussian noise and Gaussian white noise errors. The rates are controlled by the spectrum of the covariance operator of the measure that defines the error metric. Our setup is very general and allows for measures with unbounded support. A key implication is a curse of sample complexity which shows that the minimax risk for generic Lipschitz operators cannot decay at any algebraic rate in the sample size. We obtain essentially sharp characterizations when the covariance spectrum decays exponentially and provide general upper and lower bounds in slower-decay regimes.
@@ 1 Introduction
In this paper, we make a step towards closing this gap for one of the most natural and widely studied operator classes, the class of uniformly bounded Lipschitz operators. We focus on the nonparametric regression setting where the noise is either Hilbert-valued Gaussian noise (trace-class covariance operator) or Gaussian white noise. For this model class, we study the minimax risk for both fixed and random designs and quantify the decay rates as a function of the eigenvalues $\{\lambda_{i}\}_{i\geq 1}$ of the covariance operator of the measure $\mu$. Our main results give general upper and lower bounds on this quantity for classes of bounded Lipschitz operators, and we derive sharp rates in several important regimes. In particular, we remark that our setting is very general as $\mu$ could have bounded or unbounded support and includes practical cases such as the uniform measure on a compact set or an (unbounded support) Gaussian measure. This generality is in stark contrast with most works in minimax estimation.
At a qualitative level, our findings show that operator learning in the Lipschitz setting is subject to a curse of sample complexity: Regardless of how fast the eigenvalues decay, the minimax risk cannot decrease at an algebraic rate in $m$. In particular, in the case of algebraically-decaying eigenvalues, $\lambda_{i}=i^{-\alpha}$, we establish general upper and lower bounds for the minimax risk, with the former being algebraically-decaying in $\log(m)$ and the latter exponentially-decaying in $\sqrt{\log(m)}$ (note that this subalgebraic in $m$, as must be the case). For exponentially decaying eigenvalues, $\lambda_{i}=\exp(-\alpha i^{\beta})$, we obtain essentially matching upper and lower bounds for the minimax risk, leading to a precise characterization of the minimax rate in terms of $m$. Overall, these results make precise, in a minimax sense, the “curse of sample complexity” that has been identified in previous data-complexity and approximation-theoretic analyses of operator learning, which we discuss below.
@@ 1.2 Implications and related work
There has recently been a large interest in studying various statistical and data-complexity properties for operator learning problems [12, 21, 22, 32, 39, 40]. Outside of a few notable examples [29, 37], there is almost no work on minimax estimation for operator learning. The authors of [29] analyze ERM solutions over deep network classes for estimating Lipschitz operators between separable Hilbert spaces and derive nonasymptotic upper bounds for the risk. They note that sharp determination of minimax rates (and, in particular, matching lower bounds) is an open problem. On the other hand, the authors of [37] develop a general statistical theory for determining upper bounds on operators using tools from empirical-process theory. As a special case, they obtain algebraic rates for learning holomorphic operators, but note that lower bounds are unknown in the infinite-dimensional setting. In contrast, we study the minimax risk over uniformly bounded Lipschitz operators and prove information-theoretic lower bounds together with matching (or near-matching) upper bounds, implying in particular that the minimax risk cannot decay algebraically in the sample size for generic Lipschitz opertors. We find sharp rates in regimes such as exponential eigenvalue decay.
@@ 1.3 Road map and paper organization
Section 2 introduces our main results, and, in particular, states our main bounds across a range of regimes for the eigenvalue decay. The key technical aspects are the general information-theoretic lower bound proved in Section 3 and the complementary upper bound proved in Section 4. Section 5 includes the proofs of our main results. In Section 6, we present further discussion, including recovery of classical finite-dimensional minimax rates and a sharp analysis in the double-exponential eigenvalue-decay regime. We conclude in Section 7 with a number of open problems.
@@ 2 Main results and discussion
In this section, we present setup the problem, including key assumptions, then present our main results.
@@ 2.2 Main results
Suppose that Assumption 2.1 and Assumption 2.1 hold. Let $1\leq p<\infty$ and suppose that $\lambda_{i}=\exp(-\alpha i^{\beta})$ for some $\alpha>0$ and $\beta\geq 1$. Then
$\mathcal{L}_{m}(\mathcal{F}_{B,L};,L^{p}_{\mu})\asymp_{\alpha,\beta,\iota,p}\mathopen{}\mathclose{{\left(\log(m/\sigma^{2})}}\right)^{\frac{\beta}{\beta+1}},\quad m\rightarrow\infty,$
where $\mathcal{L}_{m}(\mathcal{F}_{B,L};L^{p}_{\mu})=-\log(\mathcal{M}_{m}(\mathcal{F}_{B,L};,L^{p}_{\mu}))$ is as in (2.4), with $\mathcal{M}_{m}(\mathcal{F}_{B,L};,L^{p}_{\mu})$ given by either (2.2) or (2.3).
@@
Suppose that Assumption 2.1 and Assumption 2.1 hold. Let $t>0$, $1\leq p<\infty$ and suppose that $w_{i}=i$ and $\lambda_{i}=\exp(-\alpha i^{\beta})$ for some $\alpha>0$ and $\beta\geq 1$. Then
$\mathcal{L}_{m}(\mathcal{F}^{t}_{B,L};,L^{p}_{\mu})\asymp_{\alpha,\beta,\iota,t,p}\mathopen{}\mathclose{{\left(\log(m/\sigma^{2})}}\right)^{\frac{\beta}{\beta+1}},\quad m\rightarrow\infty.$
@@
$\mathcal{M}_{m}(\mathcal{F}_{B,L};L^{p}_{\mu})\asymp C_{1}\exp\mathopen{}\mathclose{{\left(-C_{2}\mathopen{}\mathclose{{\left(\log(m/\sigma^{2})}}\right)^{\frac{\beta}{\beta+1}}}}\right),\quad m\rightarrow\infty,$
where $C_{1}=C_{1}(\alpha,\beta,\iota,\Upsilon,B,L,p)$ is a constant, potentially depending on all parameters, and $C_{2}=C_{2}(\alpha,\beta,\iota,p)$. In particular, the minimax risk is subalgebraic, but superlogarithmic, regardless of the exponential decay rate $\beta$ of the eigenvalues $\lambda_{i}$. Note that we do not specify the rate constant $C_{2}$ in this work. With additional work in tracking constants, our proofs would yield upper and lower bounds $C_{2}$. However, finding an exact expression for $C_{2}$ is an open problem.
The constants in the lower and upper bounds implied by the $\asymp_{\alpha,\beta,\iota,p}$ in Theorem 2.5 symbol may only depend on only a subset of the parameters $\alpha,\beta,\iota,p$. For succinctness, we have not made this explicit. The specific parameters appearing in the lower and upper bounds can be readily inferred from the proofs of Theorem 2.5,2.6. Note that the same consideration applies to all subsequent results in this section.
@@
Our tight characterization only holds for $\beta\geq 1$. For $0<\beta<1$, we have the non-tight bounds
$\mathopen{}\mathclose{{\left(\log(m/\sigma^{2})}}\right)^{\frac{\beta}{\beta+1}}\lesssim_{\alpha,\beta,\iota,p}\mathcal{L}_{m}(\mathcal{F}_{B,L};L^{p}_{\mu})\lesssim_{\alpha,\beta,\iota,p}\mathopen{}\mathclose{{\left(\log(m/\sigma^{2})}}\right)^{\frac{1}{2}}$
in the case of Hilbert-valued noise. An analogous result holds for white noise. In particular, the upper bound for $\mathcal{L}_{m}$ (i.e., lower bound for $\mathcal{M}_{m}$) has exponent $1/2$ independent of $\beta\in(0,1)$. We believe this is an artefact of the proof. See Remark 5.1 for further discussion.
@@
Suppose that Assumption 2.1 and Assumption 2.1 hold. Let $1\leq p<\infty$ and suppose that $\lambda_{i}=i^{-\alpha}$ for some $\alpha>1$. Then
$\frac{\alpha-1}{2}\log\mathopen{}\mathclose{{\left(\frac{\log(m/\sigma^{2})}{\log(\log(m/\sigma^{2}))}}}\right)+\mathcal{O}_{\alpha,\beta,\iota,\Upsilon,B,L,p}(1)\leq\mathcal{L}_{m}(\mathcal{F}_{B,L};L^{p}_{\mu})\lesssim_{\iota,p}\sqrt{\log(m/\sigma^{2})},$
@@
Suppose that Assumption 2.1 and Assumption 2.1 hold. Let $1\leq p<\infty$ and suppose that $\lambda_{i}=i^{-\alpha}$ for some $\alpha>1$. Then
$\frac{\alpha-1}{2}\mathopen{}\mathclose{{\left(\frac{\log(m/\sigma^{2})}{\log(\log(m/\sigma^{2}))}}}\right)+\mathcal{O}_{\alpha,\beta,\iota,t,B,L,p}(1)\leq\mathcal{L}_{m}(\mathcal{F}^{t}_{B,L};L^{p}_{\mu})\lesssim_{\iota,p}\sqrt{\log(m/\sigma^{2})}.$
@@
$C_{1}\exp\mathopen{}\mathclose{{\left(-C_{2}\sqrt{\log(m/\sigma^{2})}}}\right)\leq\mathcal{M}_{m}(\mathcal{F};L^{p}_{\mu})\leq C_{3}\mathopen{}\mathclose{{\left(\frac{\log(m/\sigma^{2})}{\log(\log(m/\sigma^{2}))}}}\right)^{-\frac{\alpha-1}{2}}$
for suitable constants $C_{1},C_{2},C_{3}>0$ depending on the parameters implied by the theorems. The lower bound is independent of algebraic decay rate $\alpha$ of the eigenvalues—notably, it is the same rate as holds in the case of exponentially-decaying eigenvalues with $\beta=1$—and therefore we suspect it is nonsharp. In fact, we believe the minimax rate behaves like
$\mathopen{}\mathclose{{\left(\log(m/\sigma^{2})}}\right)^{-\frac{\alpha}{2}},\quad m\rightarrow\infty.$
@@
To summarize, in this section we have shown that minimax rates for Lipschitz operators are necessarily subalgebraic, and have giving a series of upper and lower bounds for different eigenvalues decay profiles. We remark in passing that, while impossible, rates that are close to algebraic can be obtained for sufficiently-fast decaying eigenvalues. In Section 6.2 we present a tight characterization for the log-minimax risk in the case of double-exponential decay, i.e., $\lambda_{i}=\exp(-\exp(\alpha i))$. This corresponding minimax rate is algebraic for exponentially-large regimes of $m/\sigma^{2}$.
@@ 3 Lower bounds on the minimax risk
$\frac{1}{M}\sum_{j=1}^{M}D(P_{j}\|P_{0})\leq\alpha\log M,\quad 0<\alpha<1/8,$ (3.5)
then no estimator can reliably distinguish the hypotheses. In which case, Fano’s inequality (reproduced below) provides a uniform lower bound of order $s^{*}$ for the minimax risk.
It remains to construct the well-separated set $\{F_{0},\dots,F_{M}\}\subset\mathcal{F}$. We do this by combining (i) a geometric construction of many localized “bump” functions in the first $d$ eigencoordinates of $\mu$ with (ii) the Varshamov–Gilbert bound (also reproduced below) to extract a subset of $\{0,1\}^{n}$ with large Hamming separation, which translates into $L^{p}_{\mu}$-separation of the corresponding operators. We now recall the two standard tools (Fano’s inequality and the Varshamov–Gilbert bound) that formalize this recipe.
@@
    $\frac{1}{M}\sum_{j=1}^{M}D(P_{j}\|P_{0})\leq\alpha\log M$, with $0<\alpha<1/8$.
@@
$\inf_{\widehat{F}}\sup_{F\in\mathcal{F}}\mathbb{P}\mathopen{}\mathclose{{\left[d(\widehat{F},F)\geq s^{*}}}\right]\geq\frac{\sqrt{M}}{1+\sqrt{M}}\mathopen{}\mathclose{{\left(1-2\alpha-2\sqrt{\frac{\alpha}{\log M}}}}\right)>0$ (3.6)
@@
$\inf_{\widehat{F}}\sup_{F\in\mathcal{F}}\mathbb{E}\mathopen{}\mathclose{{\left[d(\widehat{F},F)}}\right]\geq\frac{\sqrt{M}}{1+\sqrt{M}}\mathopen{}\mathclose{{\left(1-2\alpha-2\sqrt{\frac{\alpha}{\log M}}}}\right)s^{*}.$ (3.7)
@@
Step 4: Choosing $n$ and $C_{1},\ldots,C_{n}$. We require that $n\geq 8$ and the sets $\mathcal{E}_{h}(C_{i})$ are disjoint and contained in $\mathcal{E}$. The latter assumption is equivalent to assuming that the sets $D_{h}(c_{i})$ are disjoint and contained in $D$. This is equivalent to the question of packing the unit $\infty$-norm ball $D=\{x\in\mathbb{R}^{d}:{\|x\|}_{\infty}\leq a\}$ in $\mathbb{R}^{d}$ with $\infty$-norm balls of radius $h$. Hence, we can choose $n$ as the packing number of $B$ with radius $h$. This packing number is bounded below by $(a/h)^{d}$. Since $h\leq a/8$ by assumption, we have $n\geq(a/h)^{d}\geq a/h\geq 8$, as required. Substituting this into (3.10) and recalling that $M\geq 2^{n/8}$ now yields the result. ∎
@@
Let $0<h/a\leq 1/8$. We combine the previous lemma with Fano’s inequality (Theorem 3.3). Consider the functionals $F_{0},\ldots,F_{M}$ from the previous lemma. We use these to construct operators by multiplying by a suitable element $Y\in\mathcal{Y}$: namely, $G_{j}(X)=YF_{j}(X)$ for $X\in\mathcal{X}$ and $j=0,\ldots,M$. The choice of $Y$ is dictated by the noise model, as follows:
@@
where the final inequality is due to (3.2). Hence, in (i) we have $G_{j}\in\mathcal{F}_{B,L}$, $\forall j$, and in (ii) we have $G_{j}\in\mathcal{F}^{t}_{B,L}$, $\forall j$. Let $d$ denote the $L^{p}_{\mu}$-distance, set $\alpha=1/16$ (this choice is arbitrary) and set
@@
Consider the fixed designs. In case (i), we have
$D(P_{j}\|P_{0})=\frac{1}{2\sigma^{2}}\sum_{i=1}^{m}{\|\Upsilon^{-1/2}G_{j}(X_{i})\|}^{2}\leq\frac{m}{2\sigma^{2}\upsilon_{1}}\max_{X\in\mathcal{X}}|F_{j}(X)|^{2}\leq c\frac{L^{2}h^{2}m}{\sigma^{2}\sum^{d}_{j=1}1/\lambda_{j}},$
@@
$\mathcal{M}_{m}(\mathcal{F};D)\gtrsim\frac{L}{\sqrt{\sum^{d}_{j=1}1/\lambda_{j}}}\mathopen{}\mathclose{{\left(\frac{\iota}{p+1}}}\right)^{d/p}h,$

### v2

@@ Abstract
We develop a minimax theory for operator learning, where the goal is to estimate an unknown operator between separable Hilbert spaces from finitely many noisy input-output samples. For uniformly bounded Lipschitz operators, we prove information-theoretic lower bounds together with matching or near-matching upper bounds, covering both fixed and random designs under Hilbert-valued Gaussian noise and Gaussian white noise errors. The rates are controlled by the spectrum of the covariance operator of the measure that defines the error metric. Our setup is very general and allows for measures with unbounded support. A key implication is a curse of sample complexity, which shows that the minimax risk for generic Lipschitz operators cannot decay at any algebraic rate in the sample size. We obtain sharp characterizations when the covariance spectrum decays exponentially and provide general upper and lower bounds in slower-decay regimes. Finally, we show that assuming higher regularity, i.e., Hölder smoothness, does not improve minimax rates over the Lipschitz case, up to potential constants. Thus, we show that learning operators of any finite regularity necessarily suffers a curse of sample complexity.
Keywords: operator learning, minimax rates, Lipschitz operators, Hölder operators, curse of sample complexity
Corresponding author: ben_adcock@sfu.ca (Ben Adcock)
@@ 1 Introduction
In this paper, we make a step towards closing this gap for one of the most natural and widely studied operator classes, the class of uniformly bounded Lipschitz operators. We focus on the nonparametric regression setting where the noise is either Hilbert-valued Gaussian noise (trace-class covariance operator) or Gaussian white noise. For this model class, we study the minimax risk for both fixed and random designs and quantify the decay rates as a function of the eigenvalues $\{\lambda_{i}\}_{i\geq 1}$ of the covariance operator of the measure $\mu$. Our main results give general upper and lower bounds on this quantity for classes of bounded Lipschitz operators, and we derive sharp rates in several important regimes. In particular, we remark that our setting is very general as $\mu$ could have bounded or unbounded support and includes practical cases such as the uniform measure on a compact set or an (unbounded support) Gaussian measure. This generality is in stark contrast with most works in minimax estimation. Finally, we show that imposing higher smoothness does not help: for operators of higher-order Hölder smoothness, the minimax rates do not improve over those of the Lipschitz case, except for potential constants.
At a qualitative level, our findings show that operator learning is subject to a curse of sample complexity: Regardless of how fast the eigenvalues decay, the minimax risk cannot decrease at an algebraic rate in $m$. In particular, in the case of algebraically-decaying eigenvalues, $\lambda_{i}=i^{-\tau}$, we establish general upper and lower bounds for the minimax risk, with the former being algebraically-decaying in $\log(m)$ and the latter exponentially-decaying in $\sqrt{\log(m)}$ (note that this subalgebraic in $m$, as must be the case). For exponentially decaying eigenvalues, $\lambda_{i}=\exp(-\tau i^{\omega})$, we obtain essentially matching upper and lower bounds for the minimax risk, leading to a precise characterization of the minimax rate in terms of $m$. Overall, these results make precise, in a minimax sense, the curse of sample complexity that has been identified in previous data-complexity and approximation-theoretic analyses of operator learning, which we discuss below.
@@ 1.2 Implications and related work
There has recently been a large interest in studying various statistical and data-complexity properties for operator learning problems [12, 21, 22, 32, 39, 40]. Outside of a few notable examples [29, 37], there is almost no work on minimax estimation for operator learning. The authors of [29] analyze ERM solutions over deep network classes for estimating Lipschitz operators between separable Hilbert spaces and derive nonasymptotic upper bounds for the risk. They note that sharp determination of minimax rates (and, in particular, matching lower bounds) is an open problem. On the other hand, the authors of [37] develop a general statistical theory for determining upper bounds on operators using tools from empirical-process theory. As a special case, they obtain algebraic rates for learning holomorphic operators, but note that lower bounds are unknown in the infinite-dimensional setting. In contrast, we study the minimax risk over uniformly bounded Lipschitz operators and smoother counterparts and prove information-theoretic lower bounds together with matching (or near-matching) upper bounds, implying in particular that the minimax risk cannot decay algebraically in the sample size for generic opertors. We find sharp rates in regimes such as exponential eigenvalue decay.
@@ 1.3 Road map and paper organization
Section 2 introduces our main results in the Lipschitz case, and, in particular, states our main bounds across a range of regimes for the eigenvalue decay. The key technical aspects are the general information-theoretic lower bound proved in Section 3 and the complementary upper bound proved in Section 4. Section 5 includes the proofs of these main results. In Section 6, we present further discussion, including the recovery of classical finite-dimensional minimax rates, a sharp analysis in the double-exponential eigenvalue-decay regime, and show that higher-order (Hölder) regularity does not improve the minimax rates. We conclude in Section 7 with a number of open problems.
@@ 2 Main results and discussion
In this section, we setup the problem, including key assumptions, then present our main results.
@@ 2.2 Main results
Suppose that Assumption 2.1 and Assumption 2.1 hold. Let $1\leq p<\infty$ and suppose that $\lambda_{i}=\exp(-\tau i^{\omega})$ for some $\tau>0$ and $\omega\geq 1$. Then
$\mathcal{L}_{m}(\mathcal{F}_{B,L};L^{p}_{\mu})\asymp_{\tau,\omega,\iota,p}\mathopen{}\mathclose{{\left(\log(m/\sigma^{2})}}\right)^{\frac{\omega}{\omega+1}},\quad m\rightarrow\infty,$
where $\mathcal{L}_{m}(\mathcal{F}_{B,L};L^{p}_{\mu})=-\log(\mathcal{M}_{m}(\mathcal{F}_{B,L};L^{p}_{\mu}))$ is as in (2.4), with $\mathcal{M}_{m}(\mathcal{F}_{B,L};L^{p}_{\mu})$ given by either (2.2) or (2.3).
@@
Suppose that Assumption 2.1 and Assumption 2.1 hold. Let $t>0$, $1\leq p<\infty$ and suppose that $w_{i}=i$ and $\lambda_{i}=\exp(-\tau i^{\omega})$ for some $\tau>0$ and $\omega\geq 1$. Then
$\mathcal{L}_{m}(\mathcal{F}^{t}_{B,L};,L^{p}_{\mu})\asymp_{\tau,\omega,\iota,t,p}\mathopen{}\mathclose{{\left(\log(m/\sigma^{2})}}\right)^{\frac{\omega}{\omega+1}},\quad m\rightarrow\infty.$
@@
$\mathcal{M}_{m}(\mathcal{F}_{B,L};L^{p}_{\mu})\asymp C_{1}\exp\mathopen{}\mathclose{{\left(-C_{2}\mathopen{}\mathclose{{\left(\log(m/\sigma^{2})}}\right)^{\frac{\omega}{\omega+1}}}}\right),\quad m\rightarrow\infty,$
where $C_{1}=C_{1}(\tau,\omega,\iota,\Upsilon,B,L,p)$ is a constant, potentially depending on all parameters, and $C_{2}=C_{2}(\tau,\omega,\iota,p)$. In particular, the minimax risk is subalgebraic, but superlogarithmic, regardless of the exponential decay rate $\omega$ of the eigenvalues $\lambda_{i}$. Note that we do not specify the rate constant $C_{2}$ in this work. With additional work in tracking constants, our proofs would yield upper and lower bounds $C_{2}$. However, finding an exact expression for $C_{2}$ is an open problem.
The constants in the lower and upper bounds implied by the $\asymp_{\tau,\omega,\iota,p}$ in Theorem 2.5 symbol may only depend on only a subset of the parameters $\tau,\omega,\iota,p$. For succinctness, we have not made this explicit. The specific parameters appearing in the lower and upper bounds can be readily inferred from the proofs of Theorem 2.5,2.6. Note that the same consideration applies to all subsequent results in this section.
@@
Our tight characterization only holds for $\omega\geq 1$. For $0<\omega<1$, we have the non-tight bounds
$\mathopen{}\mathclose{{\left(\log(m/\sigma^{2})}}\right)^{\frac{\omega}{\omega+1}}\lesssim_{\tau,\omega,\iota,p}\mathcal{L}_{m}(\mathcal{F}_{B,L};L^{p}_{\mu})\lesssim_{\tau,\omega,\iota,p}\mathopen{}\mathclose{{\left(\log(m/\sigma^{2})}}\right)^{\frac{1}{2}}$
in the case of Hilbert-valued noise. An analogous result holds for white noise. In particular, the upper bound for $\mathcal{L}_{m}$ (i.e., lower bound for $\mathcal{M}_{m}$) has exponent $1/2$ independent of $\omega\in(0,1)$. We believe this is an artefact of the proof. See Remark 5.1 for further discussion.
@@
Suppose that Assumption 2.1 and Assumption 2.1 hold. Let $1\leq p<\infty$ and suppose that $\lambda_{i}=i^{-\tau}$ for some $\tau>1$. Then
$\frac{\tau-1}{2}\log\mathopen{}\mathclose{{\left(\frac{\log(m/\sigma^{2})}{\log(\log(m/\sigma^{2}))}}}\right)+\mathcal{O}_{\tau,\omega,\iota,\Upsilon,B,L,p}(1)\leq\mathcal{L}_{m}(\mathcal{F}_{B,L};L^{p}_{\mu})\lesssim_{\iota,p}\sqrt{\log(m/\sigma^{2})},$
@@
Suppose that Assumption 2.1 and Assumption 2.1 hold. Let $1\leq p<\infty$ and suppose that $\lambda_{i}=i^{-\tau}$ for some $\tau>1$. Then
$\frac{\tau-1}{2}\mathopen{}\mathclose{{\left(\frac{\log(m/\sigma^{2})}{\log(\log(m/\sigma^{2}))}}}\right)+\mathcal{O}_{\tau,\omega,\iota,t,B,L,p}(1)\leq\mathcal{L}_{m}(\mathcal{F}^{t}_{B,L};L^{p}_{\mu})\lesssim_{\iota,p}\sqrt{\log(m/\sigma^{2})}.$
@@
$C_{1}\exp\mathopen{}\mathclose{{\left(-C_{2}\sqrt{\log(m/\sigma^{2})}}}\right)\leq\mathcal{M}_{m}(\mathcal{F};L^{p}_{\mu})\leq C_{3}\mathopen{}\mathclose{{\left(\frac{\log(m/\sigma^{2})}{\log(\log(m/\sigma^{2}))}}}\right)^{-\frac{\tau-1}{2}}$
for suitable constants $C_{1},C_{2},C_{3}>0$ depending on the parameters implied by the theorems. The lower bound is independent of algebraic decay rate $\tau$ of the eigenvalues—notably, it is the same rate as holds in the case of exponentially-decaying eigenvalues with $\omega=1$—and therefore we suspect it is nonsharp. In fact, we believe the minimax rate behaves like
$\mathopen{}\mathclose{{\left(\log(m/\sigma^{2})}}\right)^{-\frac{\tau}{2}},\quad m\rightarrow\infty.$
@@
To summarize, thus far in this section we have shown that minimax rates for Lipschitz operators are necessarily subalgebraic, and have giving a series of upper and lower bounds for different eigenvalues decay profiles. We remark in passing that, while impossible, rates that are close to algebraic can be obtained for sufficiently-fast decaying eigenvalues. In Section 6.2 we present a tight characterization for the log-minimax risk in the case of double-exponential decay, i.e., $\lambda_{i}=\exp(-\exp(\tau i))$. This corresponding minimax rate is algebraic for double exponentially-large regimes of $m/\sigma^{2}$.
The results in this section, which focus exclusively on Lipschitz operators, raise an important question: Does imposing higher regularity lead to faster, in particular algebraic, rates? In Section 6.2 we show that this is not the case. The same bounds hold for $C^{k,\alpha}$ (Hölder) operators for arbitrary $k\in\mathbb{N}_{0}$ and $0<\alpha\leq 1$, with the only potential difference being in the constant. Thus imposing higher, but still finite regularity cannot overcome the curse of sample complexity.
@@ 3 Lower bounds on the minimax risk
$\frac{1}{M}\sum_{j=1}^{M}D(P_{j}\|P_{0})\leq\tau\log M,\quad 0<\tau<1/8,$ (3.5)
then no estimator can reliably distinguish the hypotheses. In which case, Fano’s inequality provides a uniform lower bound of order $s^{*}$ for the minimax risk.
It remains to construct the well-separated set $\{F_{0},\dots,F_{M}\}\subset\mathcal{F}$. We do this by combining (i) a geometric construction of many localized “bump” functions in the first $d$ eigencoordinates of $\mu$ with (ii) the Varshamov–Gilbert bound to extract a subset of $\{0,1\}^{n}$ with large Hamming separation, which translates into $L^{p}_{\mu}$-separation of the corresponding operators. We now recall the two standard tools (Fano’s inequality and the Varshamov–Gilbert bound) that formalize this recipe.
@@
    $\frac{1}{M}\sum_{j=1}^{M}D(P_{j}\|P_{0})\leq\tau\log M$, with $0<\kappa<1/8$.
@@
$\inf_{\widehat{F}}\sup_{F\in\mathcal{F}}\mathbb{P}\mathopen{}\mathclose{{\left[d(\widehat{F},F)\geq s^{*}}}\right]\geq\frac{\sqrt{M}}{1+\sqrt{M}}\mathopen{}\mathclose{{\left(1-2\kappa-2\sqrt{\frac{\kappa}{\log M}}}}\right)>0$ (3.6)
@@
$\inf_{\widehat{F}}\sup_{F\in\mathcal{F}}\mathbb{E}\mathopen{}\mathclose{{\left[d(\widehat{F},F)}}\right]\geq\frac{\sqrt{M}}{1+\sqrt{M}}\mathopen{}\mathclose{{\left(1-2\kappa-2\sqrt{\frac{\kappa}{\log M}}}}\right)s^{*}.$ (3.7)
@@
Step 4: Choosing $n$ and $C_{1},\ldots,C_{n}$. We require that $n\geq 8$ and the sets $\mathcal{E}_{h}(C_{i})$ are disjoint and contained in $\mathcal{E}$. The latter assumption is equivalent to assuming that the sets $D_{h}(c_{i})$ are disjoint and contained in $D$. This is equivalent to the question of packing the $\infty$-norm ball $D=\{x\in\mathbb{R}^{d}:{\|x\|}_{\infty}\leq a\}$ in $\mathbb{R}^{d}$ with $\infty$-norm balls of radius $h$. Hence, we can choose $n$ as the packing number of $D$ with radius $h$. This packing number is bounded below by $(a/h)^{d}$. Since $h\leq a/8$ by assumption, we have $n\geq(a/h)^{d}\geq a/h\geq 8$, as required. Substituting this into (3.10) and recalling that $M\geq 2^{n/8}$ now yields the result. ∎
@@
Let $0<h/a\leq 1/8$. We combine the previous lemma with Fano’s inequality (Theorem 3.3). Consider the functionals $F_{0},\ldots,F_{M}$ from the previous lemma. We use these to construct operators by multiplying by a suitable element $Y\in\mathcal{Y}$: Namely, $G_{j}(X)=YF_{j}(X)$ for $X\in\mathcal{X}$ and $j=0,\ldots,M$. The choice of $Y$ is dictated by the noise model as follows:
@@
where the final inequality is due to (3.2). Hence, in (i) we have $G_{j}\in\mathcal{F}_{B,L}$, $\forall j$, and in (ii) we have $G_{j}\in\mathcal{F}^{t}_{B,L}$, $\forall j$. Let $d(\cdot,\cdot)$ denote the $L^{p}_{\mu}$-distance, set $\kappa=1/16$ (this choice is arbitrary) and set
@@
Consider the fixed designs case. In case (i), we have
$D(P_{j}\|P_{0})=\frac{1}{2\sigma^{2}}\sum_{i=1}^{m}{\|\Upsilon^{-1/2}G_{j}(X_{i})\|}^{2}_{\mathcal{Y}}\leq\frac{m}{2\sigma^{2}\upsilon_{1}}\max_{X\in\mathcal{X}}|F_{j}(X)|^{2}\leq c\frac{L^{2}h^{2}m}{\sigma^{2}\sum^{d}_{j=1}1/\lambda_{j}},$
@@
$\mathcal{M}_{m}(\mathcal{F};L^{p}_{\mu})\gtrsim\frac{L}{\sqrt{\sum^{d}_{j=1}1/\lambda_{j}}}\mathopen{}\mathclose{{\left(\frac{\iota}{p+1}}}\right)^{d/p}h,$